in a vessel at 70 C and 1 atm pressure, 65.6% of dinitrogen tetroxide gas dissociate into nitrogen dioxide gas. what is the partial pressure of each gas?
1
Expert's answer
2020-06-22T09:15:13-0400

Ans:

The dissociation reaction of N2O4 is:

N2O4(g)→2NO2(g)

For this question, the temperature and pressure are constant. If 65.6% dissociates of 1mol of N2O4, therefore, 0.656 mole will be consumed. Using the ICE table:

         N2O4(g)         →            2NO2(g)

Initial     1 mole                        0 mole

Change   −0.656 mole          1.312 mole

Equilibrium 0.344 mole            1.312 mole

From the equilibrium number of moles, we can find the mole fraction of each gas:

χNO2 = nNO2/ntotal = 0.344 mole/1.656 mole = 0.208

χN2O4 = nN2O4/ntotal = 1.312 mole/1.656 mole = 0.792

Since the pressure is constant Ptotal=1atm

χNO2 = PNO2/Ptotal ⇒ PNO2 = χNO2 × Ptotal = 0.208 × 1atm=0.208 atm (Ans.)

χN2O4 = PN2O4/Ptotal ⇒ PN2O4 = χN2O4 × Ptotal = 0.792 × 1atm=0.792 atm (Ans.)